# Geometric, topological and dynamical properties of conformally symplectic systems, normally hyperbolic invariant manifolds, and scattering maps

### v1

@@ 5.1. An example of a conformally symplectic map that is exact for one action form but not for another
$\begin{split}f^{*}\tilde{\alpha}_{\sigma}&=(\eta I+\mu+\varepsilon V^{\prime}(\theta)+\sigma)(d\theta+\eta dI+\varepsilon V^{\prime\prime}(\theta)d\theta)\\ &=\eta Id\theta+\eta^{2}IdI+\eta\varepsilon IV^{\prime\prime}(\theta)d\theta+(\mu+\sigma)d\theta+(\mu+\sigma)\eta dI\\ &+(\mu+\sigma)\varepsilon V^{\prime\prime}(\theta)d\theta+\varepsilon V^{\prime}(\theta)d\theta+\varepsilon\eta V^{\prime}(\theta)dI+\varepsilon^{2}V^{\prime}(\theta)V^{\prime\prime}(\theta)d\theta\\ &=\eta Id\theta+(\mu+\sigma)d\theta\\ &+d\left(\frac{\eta^{2}}{2}I^{2}+\varepsilon\eta IV^{\prime}(\theta)+\eta(\mu+\sigma)I+\varepsilon(\mu+\sigma)V^{\prime}(\theta)+\varepsilon V(\theta)+\varepsilon^{2}\frac{(V^{\prime}(\theta))^{2}}{2}\right).\end{split}$
@@
When $\eta=1$ and $\mu=0$ (5.1) becomes the conservative standard map. The drift parameter $\mu$ is fundamental in applications of the KAM theorem to conformally symplectic systems; one needs to adjust a drift parameter $\mu$ to find an invariant torus of preassigned frequency (see [CCdlL23]).
@@ 5.2. Symplectic systems that can not be perturbed into conformally symplectic ones
On the other hand the manifold would be $\mathcal{C}^{1}$-close to $\Lambda\sim{\mathbb{T}}^{2n}$ and hence have finite volume, which contradicts the fact that the dynamics on it is conformally symplectic.
A different argument based on algebraic topology, using that $\omega$ is not exact is obtained using Lemma 4.2 using that $\mathcal{C}^{1}$ perturbations are homotopic and, therefore, have the same $f^{\#2}$. Since $\eta$ is an eigenvalue of $f^{\#2}$, it cannot change from all the eigenvalues being $1$ to some of them being different from $1$. ∎
@@ 5.4. Degenerate forms in a manifold and no paring rules
However, considered as a dynamical system in $\mathbb{R}^{8}$, ${\mathcal{L}}_{(x_{4},y_{4})}$ is only a weak stable manifold for $(0,0,0,x_{4},0,0,0,y_{4})$. From the point of view of weak stable manifolds, integrability of the foliation is surprising (see [JPdlL95]).
@@ 6. Vanishing lemmas
This section is devoted to formulating and proving vanishing lemmas which are an important ingredient in the proofs of the main results of Section 3.
@@
This idea is very general and applies in many other contexts and other definition of rates, for example Sacker-Sell spectrum, Lyapunov exponents and for other geometries such as locally conformal systems.
@@ 6.2. General vanishing lemmas
The idea is very simple. If we assume that $\omega_{\mid L}$ is non-degenerate, for any $x\in L$, and $u\neq 0$ we have $\iota_{u}(\omega(x))\neq 0$, i.e., there exists $v$ such that $\omega(x)(u,v)\neq 0$. Hence, the hypothesis of the previous lemma have to fail. If for a point $x$ in an invariant manifold with a symplectic form there is a vector decreasing exponentially fast, there has to be another one growing exponentially with a rate that matches. This is the key to pairing rules, but one has to fix some details of uniformity.
@@
Assume that the symplectic form $\omega$ is uniformly bounded in a neighborhood of $L$, and that $\omega_{\mid L}$ is non-degenerate. Take $x\in L$ and assume that for some $0<\alpha<1$, there exists $u$ that satisfies (A1) of Lemma 6.2.
Then, for any $\beta$ with $\alpha\beta<\eta$, there exists $v$ that fails (A2) of Lemma 6.2
@@
For the experts in Fenichel theory, we point out that the uniformity lemma [Fen74] allows to go from (6.6) to bounds with uniform constants. Unfortunately, one of the assumptions of the uniformity lemma is compactness of the manifold, which is not true in our setting.
@@ 6.3. Vanishing lemmas on NHIMs
$\omega(x,0)=\begin{pmatrix}\omega_{xx}(x)&0\\ 0&0\end{pmatrix}$
@@
(6.10) $\omega(x,s)=\begin{pmatrix}\omega_{xx}(x)&0\\ 0&0\end{pmatrix}+e(x,y),\textrm{ with }\|e(x,y)\|\leq CM_{\omega}\|y\|.$
@@ 6.4. Vanishing lemmas on stable/unstable manifolds
The following lemma can be considered as an analogue of Lemma 6.5 for the stable and unstable manifolds of a NHIM. It will be used in the proof of part (B) of Theorem 3.3.
@@
We adopt the standing assumptions from Section 3.1. Let $y\in W^{s,\mathrm{loc}}_{\Lambda}$ and $x\in\Lambda$ such that $y\in W^{s,\mathrm{loc}}_{x}$.
Then, we have:
@@
Analogously, let $y\in W^{u,\mathrm{loc}}_{\Lambda}$ and $x\in\Lambda$ such that $y\in W^{u,\mathrm{loc}}_{x}$. Then, we have:
@@
(6.14) $\begin{split}&\forall y\in W^{s}_{x},v_{s}\in T_{y}W^{s}_{x}\implies i(v_{s})(\omega|_{W^{s}_{\Lambda}})=0,\\ &\forall y\in W^{u}_{x},v_{u}\in T_{y}W^{s}_{x}\implies i(v_{u})(\omega|_{W^{s}_{\Lambda}})=0.\end{split}$
In other words, $W^{s}_{\Lambda}$ is presymplectic and the foliation given by the kernel is the foliation of strong stable manifolds.
@@ 6.5. Results on isotropic and coisotropic manifolds
We note that $\omega_{\mid W^{s,u}_{\Lambda}}$ is presymplectic and its kernel $K_{x}(\omega)$ has constant rank equal to $d_{u}=d_{s}$ (see Remark 3.8).
@@ 6.6. Some properties of rates in isotropic invariant manifolds
Similar phenomena appear in the use of automatic reducibility in whiskered tori [CCdlL20].
@@ 6.7. Vanishing lemmas for derivatives of a general $2$-form $\omega$
This will be enough to give a proof of a variant of part (B) of Theorem 3.3 in Section 8.2.6 under the assumptions of Lemma 6.11. In particular, the proof in Section 8.2.6 does not assume that $\omega$ is closed and can be extended to cases studied in [WL98].
@@ 6.8. Vanishing lemmas for some examples of unbounded symplectic forms
In this section, we develop a result Lemma 6.12,which is very similar to Lemma 6.5, but which applies to unbounded symplectic forms. The system is assumed to have a compact invariant set $\mathcal{A}$ – that serves as the origin to measure distances, e.g. $\mathcal{A}$ could be a fixed point –. We also assume that the symplectic form at a point $x$ is bounded by a power $\alpha$ of the distance to $\mathcal{A}$ and that the hyperbolicity rates, $\alpha$, $\eta$ satisfy relations.
@@
A similar argument yields to the vanishing of the symplectic form in the case when one of the tangent vectors is unstable (or both are unstable). ∎
@@
Since the exterior derivative confutes with restriction to submanifolds, we have:
@@
In this section we show that the geometry imposes certain symmetries on the possible rates. In the case of symplectic maps, these symmetries (and their proofs) have been folklore but we have not been able to locate a specific reference. Here we derive the symmetries for conformally symplectic maps, and note that the proof also applies to the symplectic case. For conformally symplectic systems, there are arguments for periodic orbits and for Lyapunov exponents [DM96, WL98], but the argument here is different and is based on the vanishing lemmas.
@@
It is interesting to compare the proofs of pairing rules for rates above with the proofs of pairing rules for periodic orbits or for Lyapunov exponents in [DM96, WL98]. The proofs in the above references are based on defining the operator $J_{x}:T_{x}M\rightarrow T_{x}M$ by $\omega(x)(u,v)=g_{x}(u,J_{x}v)$ where $g$ is the Riemannian metric.
@@
The first proof, given in Section 8.2.1 is based on vanishing lemmas. A proof adapting the one from [DdlLS08] from the symplectic case to the conformally symplectic case is given in Section 8.2.2. In Section 8.2.3 we give a proof based on the system of coordinates defined in Section 2.8. In Section 8.2.4 we give a proof based on Cartan’s magic formula. These four proofs use the standing assumptions from Section 3.1 and conditions (S). They use strongly that $\omega$ is closed, but they do not use that $\omega$ is non-degenerate, so that these proofs apply to presymplectic forms.
@@
It is enough to prove (8.2) for vectors $v^{1},v^{2}\in T_{y}W^{s}_{\Lambda}$ that are transverse to the fiber $W^{s}_{\Omega_{+}(y)}$, that is, $v^{1},v^{2}\not\in T_{y}W^{s}_{\Omega_{+}(y)}$. Since the transversality condition is open and dense and $\omega$ is continuous, this implies (8.2), for all $v^{1},v^{2}\in T_{y}W^{s}_{\Lambda}$.
@@
Using that the projections $\Omega_{+}^{\Gamma}$ satisfy the equivariance relations (2.23), we can write
(8.7) $\begin{split}&\Omega_{+}^{\Gamma}=f_{\mid\Lambda}^{-n}\circ\Omega_{+}^{f^{n}(\Gamma)}\circ f^{n}.\end{split}$
These relations will be important later.
8.2.2. A proof of Theorem 3.3 (B) by adapting the proof of [DdlLS08] from the symplectic case to the conformally symplectic case
The proof of [DdlLS08] uses a similar geometric construction as the proof in Section 8.2.1. The paper [DdlLS08] starts from the same cell depicted in Fig. 4 and obtains the desired result by showing that the integral of $\omega$ over $\Upsilon$ is zero.
@@
The proof in Section 8.2.1, can be considered as a “disintegration” of the argument in [DdlLS08]. We can think of the vanishing lemma as dividing $\Upsilon$ into infinitesimal cells and showing that each infinitesimal integral is exactly zero. Proving first the infinitesimal result gives more flexibility and the vanishing lemmas are used also to prove the pairing rules.
@@
When $\omega_{\mid\Lambda}$ is non-degenerate, the kernel of $\omega|_{W^{s,\textrm{loc}}_{\Lambda}}$ in this neighborhood is the tangent to the $W^{s}_{x}$ leaves of the strong stable foliation 999 This is consistent with Lemma 3.6 that shows that the kernel of a presymplectic form integrates to a foliation..
@@
We hope that this new proof provides some insights that can be used to develop perturbation theories or to extend the theory to other contexts involving non-closed forms as in [WL98].
@@ 8.2.7.2 Systematic construction of approximations to scattering map
Even if Lemma 8.7 is not used in this paper, we could use it to control perturbations (some version of this was used in [GdlLM21]) or to generate numerical approximations.
@@
Using the estimates on the rates of growth of higher derivatives from [DdlLS08, Proposition 15] – recall that in this paper we are assuming condition (N) – and condition (U3) we obtain:
@@
  • By hypothesis 8.21, $\omega(\Upsilon^{n}\circ f^{n}(\mathcal{A}))=\eta^{n}a+\varepsilon_{n}|f^{n}(\mathcal{A})|$ with $\varepsilon_{n}\rightarrow 0$.
@@
Let $\sigma\subset W^{s,\mathrm{loc}}_{\Lambda}$ be a 1D-cell transversal to the foliation (2.15), parameterized by
@@
An important observation is that using the function $P^{+}$ in (8.23), by (8.22) we obtain:
@@
$\frac{d}{dt}\phi_{t}^{*}\alpha=\phi^{*}_{t}\left[i(V)d\alpha+d(i(V)\alpha)\right]=\phi_{t}^{*}(di(V)\alpha)=d(\phi^{*}_{t}(i(V)\alpha)),$
where we have used $i(V)(d\alpha)=i(V)(\omega)=0$. The last equality is because of the Vanishing Lemma 6.7.
@@
$((\Omega_{+}^{\Gamma})^{-1})^{*}\alpha-\alpha=\phi_{1}^{*}\alpha-\phi^{*}_{0}\alpha=d\int_{0}^{1}\phi^{*}_{t}(i(V)\alpha)\,dt,$
@@
$(\Omega_{+}^{\Gamma})^{*}\alpha-\alpha=-d\left[\int_{0}^{1}\left(\phi_{t}\circ(\Omega_{+}^{\Gamma}))^{*}(i(V)\alpha)\,dt\right)\right].$
Note that this proof also gives that if we have a presymplectic manifold and consider the foliation by the kernel, the holonomy maps between two transversals (which are symplectic manifolds since the kernel is excluded) are symplectic maps.

### v2

@@ 5.1. An example of a conformally symplectic map that is exact for one action form but not for another
$\begin{split}f^{*}&\tilde{\alpha}_{\sigma}=(\eta I+\mu+\varepsilon V^{\prime}(\theta)+\sigma)(d\theta+\eta dI+\varepsilon V^{\prime\prime}(\theta)d\theta)\\ &=\eta Id\theta+\eta^{2}IdI+\eta\varepsilon IV^{\prime\prime}(\theta)d\theta+(\mu+\sigma)d\theta+(\mu+\sigma)\eta dI\\ &+(\mu+\sigma)\varepsilon V^{\prime\prime}(\theta)d\theta+\varepsilon V^{\prime}(\theta)d\theta+\varepsilon\eta V^{\prime}(\theta)dI+\varepsilon^{2}V^{\prime}(\theta)V^{\prime\prime}(\theta)d\theta\\ &=\eta Id\theta+(\mu+\sigma)d\theta\\ &+d\left(\frac{\eta^{2}}{2}I^{2}+\varepsilon\eta IV^{\prime}(\theta)+\eta(\mu+\sigma)I+\varepsilon(\mu+\sigma)V^{\prime}(\theta)+\varepsilon V(\theta)+\varepsilon^{2}\frac{(V^{\prime}(\theta))^{2}}{2}\right).\end{split}$
@@
When $\eta=1$ and $\mu=0$ (5.1) becomes the conservative standard map. The drift parameter $\mu$ is fundamental in applications of the KAM theorem to conformally symplectic systems; one needs to adjust a drift parameter $\mu$ to find an invariant torus of preassigned frequency (see [CallejaCL13]).
@@
Note that the limit $\eta\to 1$ is singular. If we fix $\mu\neq 0$, the $\sigma_{*}(\eta)$ that makes (5.1) exact, goes to $\infty$ as $\eta$ converges to $1$.
An interesting variant of (5.1) (Fröeschlé map) is to take the variables $I,\theta$ to be $n$-dimensional and the drift parameter $\mu$ to be also an $n$-dimensional vector. In such a case, $V^{\prime}$ means the gradient. The calculations presented here show that when $\eta\neq 1$, for any $\mu$ we can find a unique $n$-vector $\sigma_{*}(\mu)$ so that the map is exact for $I\cdot d\theta+\sigma(\mu)\cdot d\theta$.
@@ 5.2. Symplectic systems that can not be perturbed into conformally symplectic ones
On the other hand the manifold would be $\mathcal{C}^{1}$-close to $\Lambda\sim{\mathbb{T}}^{2n}$ and hence have finite volume, which contradicts the fact that the dynamics on it is conformally symplectic. ∎
@@ 5.4. Degenerate forms in a manifold and no paring rules
However, considered as a dynamical system in $\mathbb{R}^{8}$, ${\mathcal{L}}_{(x_{4},y_{4})}$ is only a weak stable manifold for $(0,0,0,x_{4},0,0,0,y_{4})$. From the point of view of weak stable manifolds, integrability of the foliation is surprising (see [JiangPL1995]).
@@ 6. Vanishing lemmas
This section is devoted to formulating and proving vanishing lemmas, which are an important ingredient in the proofs of the main results of Section 3.
@@
This idea is very general and applies in many other contexts and other types of rates, for example Sacker-Sell spectrum and Lyapunov exponents, and for other geometries such as locally conformal systems.
@@ 6.2. General vanishing lemmas
If we assume that $\omega_{\mid L}$ is non-degenerate, for any $x\in L$, and $0\neq u\in T_{x}L$ we have $\iota_{u}(\omega(x))\neq 0$, i.e., there exists $v\in T_{x}L$ such that $\omega(x)(u,v)\neq 0$. Hence, the hypothesis of the previous lemma have to fail. If for a point $x$ in an invariant manifold with a symplectic form there is a vector decreasing exponentially fast, there has to be another one growing exponentially with a rate that matches. This is the key to pairing rules, but one has to fix some details of uniformity.
@@
Assume that the symplectic form $\omega$ is uniformly bounded in a neighborhood of $L$, and that $\omega_{\mid L}$ is non-degenerate. Take $x\in L$ and assume that for some $0<\alpha<1$, there exists $u\in T_{x}L$ that satisfies (A1) of Lemma 6.2.
Then, for any $\beta$ with $\alpha\beta<\eta$, there exists $v\in T_{x}L$ that fails (A2) of Lemma 6.2
@@
For the experts in Fenichel theory, we point out that the uniformity lemma [Fenichel74] allows to go from (6.6) to bounds with uniform constants. Unfortunately, one of the assumptions of the uniformity lemma is compactness of the manifold, which is not true in our setting.
@@ 6.3. Vanishing lemmas on NHIMs
$\omega(x,0)=\begin{pmatrix}\omega_{xx}(x)&0\\ 0&0\end{pmatrix}$.
@@
(6.10) $\omega(x,s)=\begin{pmatrix}\omega_{xx}(x)&0\\ 0&0\end{pmatrix}+e(x,y),\textrm{ with }\|e(x,y)\|\leq CM^{\prime}_{\omega}\|y\|.$
@@
Remark 6.7.
As a consequence of Lemma 6.5, we obtain that the stable and unstable bundles have equal dimension, i.e., $d_{s}=d_{u}$. The reason is that relative to $T_{\Lambda}M\oplus E^{s}\oplus E^{u}$ we have
$\omega=\begin{pmatrix}\omega_{xx}&0&0\\ 0&0&\omega_{y_{s}y_{u}}\\ 0&\omega_{y_{u}y_{s}}&0\\ \end{pmatrix}$
and $\omega_{y_{s}y_{u}}$, $\omega_{y_{u}y_{s}}$ are non-degenerate. Therefore, $E^{s}$ is isomorphic to $E^{u}$.
@@ 6.4. Vanishing lemmas on stable/unstable manifolds
The following lemma can be considered as an analogue of Lemma 6.5 for the stable and unstable manifolds of a NHIM. It will be used in the proof of part (B) of Theorem  3.3.
@@
We adopt the standing assumptions from Section 3.1.
Let $y\in W^{s,\mathrm{loc}}_{x}$ for some $x\in\Lambda$. Then, we have:
@@
Analogously, let $y\in W^{u,\mathrm{loc}}_{x}$ for some $x\in\Lambda$. Then, we have:
@@
Note that in (6.11), we have $\lambda_{+}\mu_{+}\eta^{-1}<1$ implies $\lambda_{+}^{2}\eta^{-1}<1$, since $\lambda_{+}<\mu_{+}$.
@@
(6.14) $\begin{split}&\forall y\in W^{s}_{x},v_{s}\in T_{y}W^{s}_{x}\implies i_{v_{s}}(\omega_{\mid W^{s}_{\Lambda}})=0,\\ &\forall y\in W^{u}_{x},v_{u}\in T_{y}W^{u}_{x}\implies i_{v_{u}}(\omega_{\mid W^{u}_{\Lambda}})=0.\end{split}$
In other words, $W^{s,u}_{\Lambda}$ is presymplectic and the foliation given by the kernel is the foliation of strong stable/unstable manifolds.
@@ 6.5. Results on isotropic and coisotropic manifolds
We note that $\omega_{\mid W^{s,u}_{\Lambda}}$ is presymplectic and its kernel $K_{x}(\omega)$ has constant rank equal to $d_{u}=d_{s}$; see Remark 3.8. Recall that $d_{s}=d_{u}$ from Remark 6.7.
@@ 6.6. Some properties of rates in isotropic invariant manifolds
Similar phenomena appear in the use of automatic reducibility in whiskered tori [CallejaCL20].
@@ 6.7. Vanishing lemmas for derivatives of a general $2$-form $\omega$
This will be enough to give a proof of a variant of part (B) of Theorem 3.3 in Section 8.2.6 under the assumptions of Lemma 6.12. In particular, the proof in Section 8.2.6 does not assume that $\omega$ is closed and can be extended to cases studied in [wojtkowski1998conformally].
@@ 6.8. Vanishing lemmas for some examples of unbounded symplectic forms
In this section, we develop a result Lemma 6.13,which is very similar to Lemma 6.5, but which applies to some unbounded symplectic forms. The system is assumed to have a compact invariant set $\mathcal{A}$ – that serves as the origin to measure distances, e.g. $\mathcal{A}$ could be a fixed point. We also assume that the symplectic form at a point $x$ is bounded by a power $\alpha$ of the distance from $x$ to $\mathcal{A}$ and that the hyperbolicity rates together with $\alpha$ and $\eta$ satisfy some relations.
@@
A similar argument gives the vanishing of the symplectic form in the case when one of the tangent vectors is unstable (or both are unstable). ∎
@@
Since the exterior derivative commutes with restriction to submanifolds, we have:
@@
In this section we show that the geometry imposes certain symmetries on the possible rates. In the case of symplectic maps, these symmetries (and their proofs) have been folklore but we have not been able to locate a specific reference. Here we derive the symmetries for conformally symplectic maps, and note that the proof also applies to the symplectic case. For conformally symplectic systems, there are arguments for periodic orbits and for Lyapunov exponents [dettmann1996proof, wojtkowski1998conformally], but the argument here is different and is based on the vanishing lemmas.
@@
It is interesting to compare the proofs of pairing rules for rates above with the proofs of pairing rules for periodic orbits or for Lyapunov exponents in [Dessler88, dettmann1996proof, wojtkowski1998conformally]. The proofs in the above references are based on defining the operator $J_{x}:T_{x}M\rightarrow T_{x}M$ by $\omega(x)(u,v)=g_{x}(u,J_{x}v)$ where $g$ is the Riemannian metric.
@@
The first proof, given in Section 8.2.1 is based on vanishing lemmas. A proof adapting the one from [DLS08] from the symplectic case to the conformally symplectic case is given in Section 8.2.2. In Section 8.2.3 we give a proof based on the system of coordinates defined in Section 2.8. In Section 8.2.4 we give a proof based on Cartan’s magic formula. These four proofs use the standing assumptions from Section 3.1 and conditions (S). They use strongly that $\omega$ is closed, but they do not use that $\omega$ is non-degenerate, so that these proofs apply to presymplectic forms.
@@
8.2.2. A proof of Theorem 3.3 (B) by adapting the proof of [DLS08] from the symplectic case to the conformally symplectic case
The proof of [DLS08] uses a similar geometric construction as the proof in Section 8.2.1. The paper [DLS08] starts from the same cell depicted in Fig. 4 and obtains the desired result by showing that the integral of $\omega$ over $\Upsilon$ is zero.
@@
The proof in Section 8.2.1, can be considered as a “disintegration” of the argument in [DLS08]. We can think of the vanishing lemma as dividing $\Upsilon$ into infinitesimal cells and showing that each infinitesimal integral is exactly zero. Proving first the infinitesimal result gives more flexibility and the vanishing lemmas are used also to prove the pairing rules.
@@
When $\omega_{\mid\Lambda}$ is non-degenerate, the kernel of $\omega|_{W^{s,\textrm{loc}}_{\Lambda}}$ in this neighborhood is the tangent to the $W^{s}_{x}$ leaves of the strong stable foliation  [^11] This is consistent with Lemma 3.6 that shows that the kernel of a presymplectic form integrates to a foliation..
@@
We hope that this new proof provides some insights that can be used to develop perturbation theories or to extend the theory to other contexts involving non-closed forms as in [wojtkowski1998conformally].
@@ 8.2.7.2 Systematic construction of approximations to scattering map
Even if Lemma 8.7 is not used in this paper, we could use it to control perturbations (some version of this was used in [GideaLM21]) or to generate numerical approximations.
@@
Using the estimates on the rates of growth of higher derivatives from [DLS08, Proposition 15] – recall that in this paper we are assuming condition (N) – and condition (U3) we obtain:
@@
  • By hypothesis 8.20, $\omega(\Upsilon^{n}\circ f^{n}(\mathcal{A}))=\eta^{n}a+\varepsilon_{n}|f^{n}(\mathcal{A})|$ with $\varepsilon_{n}\rightarrow 0$.
@@
Let $\sigma\subset W^{s,\mathrm{loc}}_{\Lambda}$ be a 1D-cell parameterized by
@@
Using the function $P^{+}$ in (8.22), by (8.21) we obtain:
@@
$\frac{d}{dt}\phi_{t}^{*}\alpha=\phi^{*}_{t}\left[i_{V}d\alpha+d(i_{V}\alpha)\right]=\phi_{t}^{*}(di_{V}\alpha)=d(\phi^{*}_{t}(i_{V}\alpha)),$
where we have used $i_{V}(d\alpha)=i_{V}(\omega)=0$. The last equality is because of the Vanishing Lemma 6.8.
@@
$((\Omega_{+}^{\Gamma})^{-1})^{*}\alpha-\alpha=\phi_{1}^{*}\alpha-\phi^{*}_{0}\alpha=d\int_{0}^{1}\phi^{*}_{t}(i_{V}\alpha)\,dt,$
@@
$(\Omega_{+}^{\Gamma})^{*}\alpha-\alpha=-d\left[\int_{0}^{1}\left(\phi_{t}\circ(\Omega_{+}^{\Gamma}))^{*}(i_{V}\alpha)\,dt\right)\right].$
Remark 8.13.
Note that this proof also gives that, if we have a presymplectic manifold and consider the foliation by the kernel, the holonomy maps between two transversals (which are symplectic manifolds since the kernel is excluded) are symplectic maps.